solve
x^2 + (x+7)^2 = 169

Answer by nerdybill(7384)

x^2 + (x+7)^2 = 169
x^2 + (x+7)(x+7) = 169
x^2 + x^2+14x+49 = 169
2x^2+14x+49 = 169
2x^2+14x-120 = 0
x^2+7x-60 = 0
(x+12)(x-5) = 0
x = {-12, 5}